Historical data indicates that only 20% of cable customers are willing to switch companies. If a binomial process is assumed, then in a sample of 20 cable customers, what is the probability that no more than 3 customers would be willing to switch their cable? a. 0.20 b. 0.15 c. 0.85 d. 0.411
Answer #1

Solution

Given that ,

Using binomial probability formula ,

P(X = x) = (n C x) * px * (1 - p)n - x

P(X < 3) = P(X = 0) + P(X = 1) + P(X = 2)

= (20 C 0) * 0.200 * (0.80)20 + (20 C 1) * 0.201 * (0.80)19 + (20 C 2) * 0.202 * (0.80)18

= 0.411

Probability = 0.411

d)